If cosec theta + cot theta = p prove that cos theta = p2-1 / p2 + 1

Grade:10

2 Answers

Yash Patari
2 years ago
cosec theta + cot theta = P (given) we know, cosec² theta - cot² theta = 1 (cosec theta + cot theta)( cosec theta - cot theta)= 1 (cosec theta - cot theta) × P = 1 ( cosec theta - cot theta) = 1/P Let's solve (i ) and (ii) cosec theta = 1+ P²/2P cot theta = P²-1/2P cot theta / cosec theta= P²-1/2P×2P/P2²+1 cos theta = P²-1/P²+1
Pasty Dawe
one year ago
God is wonderful. It still feels incredible that my life has changed. I thank Dr Amber everyday and I’m hugely grateful for his help. It is the best feeling in the world knowing that I am financially secure. Dr Amber prayed and prepared a lottery spell for me and gave me some numbers to play the lottery game. I played the set for life lottery game as Dr Amber instructed with my full hopes and trust on him to win it because I have never been known for winning anything. I was shocked, surprised and couldn’t believe what my eyes were seeing when I checked my ticket at the shop I bought it from. I was declared the winner of the set for life lottery of $67,500,000! I stopped working as a midwife after I won and ever since then, my life and that of my family has changed and I want to say a very big thank you to Dr Amber for making a tremendous change in my life. WA Dr Amber for help via +12096910783 or call +13183065044 For more enquiries or you email: amberlottotemple@ yahoo.com